Costs per Equivalent Unit

The following information concerns production in the Baking Department for December. All direct materials are placed in process at the beginning of production.

ACCOUNT Work in Process—Baking Department ACCOUNT NO.
Date Item Debit Credit Balance
Debit Credit
Dec. 1 Bal., 7,500 units, 2/5 completed 15,750
31 Direct materials, 135,000 units 243,000 258,750
31 Direct labor 70,840 329,590
31 Factory overhead 39,848 369,438
31 Goods finished, 136,800 units 355,530 13,908
31 Bal., ? units, 4/5 completed 13,908

a. Based on the above data, determine each cost listed below. Round "cost per equivalent unit" answers to the nearest cent.

1. Direct materials cost per equivalent unit

2. Conversion cost per equivalent unit

3. Cost of the beginning work in process completed during December

4. Cost of units started and completed during December

5. Cost of the ending work in process
Answer #1

a) Equivalent unit of material = 135000

Cost per equivalent unit of material = 243000/135000 = 1.8

b)Equivalent unit of conversion = (7500*3/5+129300+5700*4/5) = 138360

Cost per equivalent unit of conversion = (70840+39848)/138360 = 0.80

c) Cost of beginning WIP = 7500*3/5*.8+15750 = 19350

d) Cost of unit started and completed = 2.6*129300 = 336180

e) Cost of ending work in process = (5700*1.8+4560*.80) = 13908